OpenStudy (anonymous):

Help pleaseeeeeeeeee!!! My brain needs your help!! The average (arithmetic mean) of five numbers is 3x + 4. If one of the numbers is 7x - 4, what is the average of the other four numbers?

OpenStudy (anonymous):

I know the answer but I need to know why its 15x+20-(7x-4) rather than 15x+20-7x-4

OpenStudy (campbell_st):

Average = sum of scores/number of scores you know the average and the number 3x + 4 = sum/5 so you need to 1st find the sum of the 5 numbers, so multiply both sides by 5 5(3x + 4) = sum of the scores 15x + 20 = sum of 5 numbers and then subtract 7x - 4 to find the sum of 4 numbers sum of 4 numbers = 15x + 20 -(7x - 4) hope that makes sense

OpenStudy (anonymous):

yeah so you are smart @yomamabf

OpenStudy (anonymous):

but can you answer my question above? why does it have to be inside a parentheses rather than not?

OpenStudy (anonymous):

because you times those

OpenStudy (campbell_st):

well remember 1 number is 7x - 4 so subtracting the whole of the number is 15x + 20 - (7x - 4) = 15x + 20 - 7x + 4 after you distribute the -1 its like saying 10 minus negative 9 = 10 - (-9) = 10 + 9 = 19

OpenStudy (anonymous):

ohhhh okay because it's one number gotchaaaa thanks!!!